# Beyond Force Metrics: Pre-Training MLFFs for Stable MD Simulations

### v1

@@ Abstract
Machine-learning force fields (MLFFs) have emerged as a promising solution for speeding up ab initio molecular dynamics (MD) simulations, where accurate force predictions are critical but often computationally expensive. In this work, we employ GemNet-T, a graph neural network model, as an MLFF and investigate two training strategies: (1) direct training on MD17 (10K samples) without pre-training, and (2) pre-training on the large-scale OC20 dataset followed by fine-tuning on MD17 (10K). While both approaches achieve low force mean absolute errors (MAEs), reaching 5 meV/Å per atom, we find that lower force errors do not necessarily guarantee stable MD simulations. Notably, the pre-trained GemNet-T model yields significantly improved simulation stability, sustaining trajectories up to three times longer than the model trained from scratch. These findings underscore the value of pre-training on large, diverse datasets to capture complex molecular interactions and highlight that force MAE alone is not always a sufficient metric of MD simulation stability.
@@ 1 Introduction
Molecular dynamics (MD) simulation is a computational method used to analyze the physical movements of atoms and molecules over time, enabling the exploration of molecular interactions, structural rearrangements, and thermodynamic properties at the atomic scale. The primary aim of MD simulations is to provide insights into atomistic processes that are challenging or impossible to observe experimentally. Typical examples include folding pathways of small proteins and peptides 1, ligand binding kinetics in drug discovery campaigns 2, ion transport in solid-state battery electrolytes 3, crack propagation in structural alloys 4, and nucleation events in super-cooled liquids 5.
@@
Density functional theory (DFT) plays a critical role in molecular dynamics (MD) simulations by providing quantum mechanical calculations of interatomic forces derived from the electronic structure. While DFT yields highly accurate forces by explicitly accounting for electron interactions, its computational cost scales poorly with system size and simulation time, limiting its applicability to large-scale or long-timescale simulations 6. This limitation renders DFT impractical for long MD trajectories involving hundreds of thousands of time-steps or for simulations of systems with thousands of atoms. Consequently, classical MD simulations often rely on empirical force fields that approximate these interactions, sacrificing some accuracy to achieve feasible simulation durations and system sizes. 7 8
@@
To address the computational limitations of ab initio methods like DFT, MLFFs have emerged as promising alternatives for MD simulations 6. MLFFs are data-driven models trained to approximate the potential energy surface (PES) and interatomic forces of atomistic systems based on reference calculations from quantum mechanical methods like DFT. By learning the mapping from atomic configurations to energies and forces, MLFFs enable simulations that approach quantum-level accuracy while being orders of magnitude faster than traditional ab initio approaches. As a result, they allow the exploration of larger systems and longer timescales that would otherwise be intractable using direct quantum mechanical methods.
@@
Significant advancements have been made in the field of MLFFs, primarily through graph neural network (GNN) models such as SchNet 9, DimeNet 10, PaiNN 11, DeepPot-SE 12, NequIP 13, and GemNet 14. These models leverage graph-based representations where atoms are treated as nodes and atomic interactions as edges. SchNet introduced continuous convolutional filters to model atomic interactions, while DimeNet extended these ideas by explicitly incorporating angular information. PaiNN further refined the approach by utilizing equivariant message-passing layers that respect rotational symmetries. GemNet advanced this by incorporating directional embeddings to better capture radial and angular geometric features, providing superior handling of directional information crucial for accurate molecular force prediction.
@@
Despite their successes, MLFFs exhibit notable shortcomings, particularly regarding simulation stability and extrapolation behavior. A benchmark study conducted by Fu et al.15 emphasizes the limitations of force prediction metrics alone, demonstrating that high accuracy in force prediction does not always correlate with stable, realistic MD trajectories. They observed that models frequently failed simulations due to instability caused by extrapolation errors, where models predicted extreme, non-physical forces for configurations inadequately represented in training datasets. This instability often resulted in catastrophic trajectory divergence, bond breakage, and unrealistic configurations, even if models initially exhibited low force prediction errors. Fu et al.15 highlighted that these failures predominantly occur in simulations requiring long timescales or involving systems with complex interactions not well captured by the training data. These findings highlight stability and comprehensive configurational space coverage as critical metrics, suggesting that the development of robust MLFFs requires training on diverse molecular configurations to improve both simulation stability and accuracy.
@@
Motivated by these limitations, we explored pre-training strategies for improving MLFF performance and stability. Prior studies have shown substantial benefits of pre-training, where initial training on large and diverse datasets enhances the transferability and robustness of MLFFs. For example, pre-training has been shown to significantly enhance the generalizability and stability of models, allowing them to better capture complex molecular interactions encountered in extended simulations 16, 17, 18, 19. Additionally, recent work has shown that pretrained models, such as EquiformerV2, can be fine-tuned to achieve state-of-the-art performance even on challenging, out-of-domain systems like high-entropy alloys, significantly accelerating atomistic discovery in materials science 20. More broadly, various forms of pre-training allow graph neural networks to develop transferable atomic representations across diverse domains. These include small molecules, catalysts, and inorganic crystals, fostering more universal chemical understanding 21, 22, 23.
@@
In this study, we investigate how pre-training an MLFF, specifically GemNet-T, on large-scale datasets impacts the stability of MD simulations. We show that pre-training on diverse atomic systems, including catalyst datasets, followed by fine-tuning on smaller molecular datasets, effectively reduces overfitting and enhances generalizability. These results underscore the significance of comprehensive, large-scale pre-training in capturing complex molecular interactions and provide a foundational framework for developing robust MLFFs for MD simulations.
@@ 2.1 Molecular Dynamics Simulation
MD simulation is a computational method used to model the physical movements of atoms and molecules over time by solving Newton’s equations of motion24. The MD simulation follows the Newtonian equation of motion, given by:
@@
To control the temperature, a Nosé–Hoover thermostat was employed to maintain a target temperature of 500 K 15. The simulation was conducted using a time step of 0.5 fs and was run for 600,000 steps, corresponding to 300 ps of dynamics. Atomic positions were periodically saved for subsequent analysis.
@@ 2.2 Structural Validation
To assess the structural fidelity of MLFFs in MD simulations of small molecules like Aspirin, we compute the pair-distance distribution function, $h(r)$. Unlike the radial distribution function (RDF) used in bulk or periodic systems, the pair-distance distribution function is better suited for finite molecular systems 15, capturing the distribution of all interatomic distances within a molecule. It is formally defined as:
@@
where $N$ is the number of atoms and $\mathbf{x}_{i}$ are their positions. This function provides a low-dimensional, geometry-sensitive descriptor of molecular structure, enabling direct comparisons between model-predicted and reference (ab initio) simulations. Following prior work 17, we are also able to compute the mean absolute error (MAE) between predicted and reference pair-distance distribution function curves to quantify structural deviations:
[…]
This metric helps detect unphysical distortions like bond stretching or angle collapse, common causes of failure in MLFF driven MD, providing a more interpretable and structure specific complement to conventional force MAE.
@@ 2.3 Stability Criterion
We assess the stability of MD simulations for flexible molecules by monitoring bond lengths over time. A simulation is classified as unstable if, at any time $t<300\,\text{ps}$, the bond length between any atom pair exceeds a specified threshold, indicating bond breaking or excessive stretching that compromises structural integrity before the intended simulation duration. The criterion for instability is expressed as 15:
@@ 2.4 Machine Learning Force Fields
MLFFs learn an approximate representation of the potential energy surface from reference data, typically generated by quantum mechanical methods such as DFT. At each simulation step, the current atomic positions are input to the MLFF, which predicts the forces acting on each atom. These predicted forces are then used to iteratively update atomic positions throughout the simulation. Recently, GNN architectures have gained prominence for effectively capturing atomic connectivity in molecular systems. GNNs capture the connectivity of atomic structures by representing 3D atomic systems as graphs, where atoms are treated as nodes and interatomic distances define the edges 14.
In this study, we utilize GemNet-T as the MLFF model. GemNet-T is a high-performing GNN designed to learn molecular energies and forces, building on prior models like DimeNet++ 14. It enhances the representation of multi-scale geometric interactions by incorporating directional embeddings, capturing both radial and angular features. This enables more accurate force predictions, particularly for complex three-dimensional atomic arrangements.
Figure 1c illustrates the core structure of the GemNet-T model, a directional message-passing neural network designed for MD simulations. The model begins by embedding atomic information using both atomic numbers and local geometric features, including distances, angles, and dihedral angles. These inputs generate initial atom and edge embeddings that preserve the molecular structure. The embeddings are then processed through multiple stacked interaction blocks, each consisting of a directional message-passing module and an atom-wise update. In GemNet-T, the message passing operates over one-hop neighborhoods and uses triplet-based geometric information (from distance and angle inputs) to encode directional relationships. This approach avoids the complexity of higher-order interactions, such as the quadruplet-based dihedral terms used in GemNet-Q, while still capturing essential directional features 14. After each message-passing step, atomic embeddings are updated through residual connections. Once all interaction layers are completed, the final atomic embeddings are aggregated to predict molecular energies and atomic forces. Importantly, these predictions are designed to be equivariant to translations, rotations, and reflections, ensuring that the model produces physically consistent results during MD simulations 14.
@@ 2.5 Datasets
We use the MD17 dataset as the primary benchmark for small-molecule systems. MD17 consists of ab initio MD trajectories, where each snapshot provides atomic coordinates, DFT total energy, and per-atom forces 25. Aspirin, a medium-sized organic molecule with 21 atoms and multiple rotatable bonds, is selected as a representative system due to its structural flexibility. This flexibility amplifies the impact of small force prediction errors, potentially leading to unstable trajectories over time. The aspirin subset in MD17 includes over 20,000 configurations (i.e., molecular snapshots along an MD trajectory, each with atomic positions, energy, and forces). For this study, we use 9,500 configurations for training, 500 for validation, and 10,000 as a separate test set.
To extend the model’s generalization beyond isolated small molecules, we incorporate the Open Catalyst 2020 (OC20) dataset for pre-training. OC20 contains 1.3 million relaxations of molecular adsorptions on surfaces, supported by over 260 million corresponding DFT calculations 26. We specifically utilize a 2-million-sample subset from the Structure-to-Energy-and-Force (S2EF) task, which provides forces and energies for each relaxed configuration. Unlike MD17, OC20 data include periodic boundary conditions, more accurately reflecting realistic catalytic surfaces. By training on this larger, chemically diverse dataset, we aim to equip the model with robust atomic interaction representations before fine-tuning on the smaller MD17 aspirin set.
@@ 3 Experiments
Figure 1: Overview of the MLFF framework for molecular dynamics (MD) simulations. a. GemNet-T is trained on molecular (MD17) and catalyst (OC20) datasets, then used as a force predictor in MD simulations. b. Atomic positions are iteratively updated based on forces predicted by the trained GemNet-T model. c. Simplified illustration of the GemNet-T architecture, which incorporates multi-hop geometric message passing and directional embeddings to model atomic interactions.
@@ 4.1 Force Metric Limitations
To comprehensively assess the impact of pre-training, we compare force prediction accuracy and MD simulation stability. The MAE in force prediction serves as the primary error metric during MLFF training, as it is the objective function the model is optimized for. To evaluate dynamical stability, we use the instability onset time ($t_{\mathrm{inst}}$), which measures how long the simulation preserves structural integrity before any bond deviates beyond a defined threshold from its reference length. A higher instability onset time, approaching the 300 ps reference from ab initio MD, indicates greater stability and closer agreement with the ground truth trajectory. The instability onset times shown in Figure 2b represent the average values obtained from three independent simulations initialized with different random seeds.
[…]
In contrast, the MD stability trends diverge from the force MAE curve. Although the force MAE clearly decreases with further training, MD stability does not necessarily improve in tandem. In some cases, instability occurs earlier at later training epochs, reflected by a reduced instability onset time, despite better force prediction accuracy. This trend is particularly pronounced in the non-pre-trained case, suggesting that the model may overfit to the training dataset and struggle to generalize to diverse configurations encountered during extended MD simulations. This observation aligns with the findings of Fu et al. 15, indicating that force metrics alone may not reliably reflect downstream performance in MD simulations.
Figure 2: Impact of pretraining on force accuracy and simulation stability. a. Force MAE (meV/Å) over epochs for the GemNet-T model trained without pretraining (MD17 only) versus with pre-training (MD17 pre-training followed by OC20 finetuning). b. Instability onset time over epochs. Error bars represent results from three independent simulations.
4.2 Stability Enhancement via Pre-Training

### v2

@@ Abstract
Machine-learning force fields (MLFFs) have emerged as a promising solution for speeding up ab initio molecular dynamics (MD) simulations, where accurate force predictions are critical but often computationally expensive. In this work, we employ GemNet-T, a graph neural network model, as an MLFF and investigate two training strategies: (1) direct training on MD17 (10K samples) without pre-training, and (2) pre-training on the large-scale OC20 dataset followed by fine-tuning on MD17 (10K). While both approaches achieve low force mean absolute errors (MAEs), reaching 5 meV/Å per atom, we find that lower force errors do not necessarily guarantee stable MD simulations. Notably, the pre-trained GemNet-T model yields significantly improved simulation stability, sustaining trajectories up to three times longer than the model trained from scratch. By analyzing local properties of the learned force fields, we find that pre-training produces more structured latent representations, smoother force responses to local geometric changes, and more consistent force differences between nearby configurations, all of which contribute to more stable and reliable MD simulations. These findings underscore the value of pre-training on large, diverse datasets to capture complex molecular interactions and highlight that force MAE alone is not always a sufficient metric of MD simulation stability.
@@ 1 Introduction
Molecular dynamics (MD) simulation is a computational method used to analyze the physical movements of atoms and molecules over time, enabling the exploration of molecular interactions, structural rearrangements, and thermodynamic properties at the atomic scale. The primary aim of MD simulations is to provide insights into atomistic processes that are challenging or impossible to observe experimentally. Typical examples include folding pathways of small proteins and peptides Lindorff-Larsen2011, ligand binding kinetics in drug discovery campaigns nunesalves2020binding, ion transport in solid-state battery electrolytes mabuchi2021ion, crack propagation in structural alloys lee2023crack, and nucleation events in super-cooled liquids sosso2016nucleation.
@@
Density functional theory (DFT) plays a critical role in MD simulations by providing quantum mechanical calculations of interatomic forces derived from the electronic structure. While DFT yields highly accurate forces by explicitly accounting for electron interactions, its computational cost scales poorly with system size and simulation time, limiting its applicability to large-scale or long-timescale simulations unke2021mlff. This limitation renders DFT impractical for long MD trajectories involving hundreds of thousands of time-steps or for simulations of systems with thousands of atoms. For this reason, classical MD simulations often rely on empirical force fields that approximate these interactions, sacrificing some accuracy to achieve feasible simulation durations and system sizes smith2018ani1, gunceler2023improving.
@@
To address the computational limitations of ab initio methods like DFT, MLFFs have emerged as promising alternatives for MD simulations unke2021mlff. MLFFs are data-driven models trained to approximate the potential energy surface (PES) and interatomic forces of atomistic systems based on reference calculations from quantum mechanical methods like DFT. By learning the mapping from atomic configurations to energies and forces, MLFFs enable simulations that approach quantum-level accuracy while being orders of magnitude faster than traditional ab initio approaches. As a result, they allow the exploration of larger systems and longer timescales that would otherwise be intractable using direct quantum mechanical methods.
@@
Significant advancements have been made in the field of MLFFs, primarily through graph neural network (GNN) models such as SchNet schutt2017schnet, DimeNet klicpera2020dimenet, PaiNN schutt2021painn, DeepPot-SE zhang2018deeppot, NequIP batzner2022nequip, and GemNet gasteiger2021gemnet. These models leverage graph-based representations where atoms are treated as nodes and atomic interactions as edges. SchNet introduced continuous convolutional filters to model atomic interactions, while DimeNet extended these ideas by explicitly incorporating angular information. PaiNN further refined the approach by utilizing equivariant message-passing layers that respect rotational symmetries. GemNet advanced this by incorporating directional embeddings to better capture radial and angular geometric features, providing superior handling of directional information crucial for accurate molecular force prediction.
@@
Despite their successes, MLFFs exhibit notable shortcomings, particularly regarding simulation stability and extrapolation behavior. A benchmark study conducted by Fu et al.fu2023forces emphasizes the limitations of force prediction metrics alone, demonstrating that high accuracy in force prediction does not always correlate with stable and realistic MD trajectories. They observed that models frequently failed simulations due to instability caused by extrapolation errors, where models predicted extreme, non-physical forces for configurations inadequately represented in training datasets. This instability often resulted in catastrophic trajectory divergence, bond breakage, and unrealistic configurations, even if models initially exhibited low force prediction errors. Fu et al.fu2023forces highlighted that these failures predominantly occur in simulations requiring long timescales or involving systems with complex interactions not well captured by the training data. These findings highlight stability and comprehensive configurational space coverage as critical metrics, suggesting that the development of robust MLFFs requires training on diverse molecular configurations to improve both simulation stability and accuracy.
@@
Motivated by these limitations, we explored pre-training strategies for improving MLFF performance and stability, with the goal of learning better molecular representations and more accurate correlations among similar structures. Prior studies have shown substantial benefits of pre-training, where initial training on large and diverse datasets enhances the transferability and robustness of MLFFs. For example, pre-training has been shown to significantly enhance the generalizability and stability of models, allowing them to better capture complex molecular interactions encountered in extended simulations chanussot2021oc20, raja2025stabilityawaretrainingmachinelearning, wang2025pfdautomaticallygeneratingmachine, peptidebert. Additionally, recent work has shown that pretrained models, such as EquiformerV2, can be fine-tuned to achieve state-of-the-art performance even on challenging, out-of-domain systems like high-entropy alloys, significantly accelerating atomistic discovery in materials science clausen2024. More broadly, various forms of pre-training allow graph neural networks to develop transferable atomic representations across diverse domains. These include small molecules, catalysts, and inorganic crystals, fostering more universal chemical understanding Wang2022, Magar2022, ock2024multimodal.
@@
In this study, we investigate how pre-training an MLFF, specifically GemNet-T, on large-scale datasets impacts the stability of MD simulations. We show that pre-training on diverse atomic systems, including catalyst datasets, followed by fine-tuning on smaller molecular datasets, effectively reduces overfitting and enhances generalizability. Pre-training yields more structured latent representations, smoother force responses to local geometric perturbations, and more consistent force differences between nearby configurations. As a result, model stability is substantially improved, even though the force prediction accuracy itself does not change significantly. These results underscore the significance of comprehensive, large-scale pre-training in capturing complex molecular interactions and provide a foundational framework for developing robust MLFFs for MD simulations.
@@ 2.1 Molecular Dynamics Simulation
MD simulation is a computational method used to model the physical movements of atoms and molecules over time by solving Newton’s equations of motionfrenkel2001understanding. The MD simulation follows the Newtonian equation of motion, given by:
@@
To control the temperature, a Nosé–Hoover thermostat was employed to maintain a target temperature of 500 K fu2023forces. The simulation was conducted using a time step of 0.5 fs and was run for 600,000 steps, corresponding to 300 ps of dynamics. Atomic positions were periodically saved for subsequent analysis.
@@ 2.2 Structural Validation
To assess the structural fidelity of MLFFs in MD simulations of small molecules like Aspirin, we compute the pair-distance distribution function, $h(r)$. Unlike the radial distribution function (RDF) used in bulk or periodic systems, the pair-distance distribution function is better suited for finite molecular systems fu2023forces, capturing the distribution of all interatomic distances within a molecule. It is formally defined as:
@@
where $N$ is the number of atoms and $\mathbf{x}_{i}$ are their positions. This function provides a low-dimensional, geometry-sensitive descriptor of molecular structure, enabling direct comparisons between model-predicted and reference (ab initio) simulations. Following prior work raja2025stabilityawaretrainingmachinelearning, we are also able to compute the MAE between predicted and reference pair-distance distribution function curves to quantify structural deviations:
$\text{MAE}_{h(r)}=\int_{0}^{\infty}\left|\langle h(r)\rangle-\langle\hat{h}(r)\rangle\right|dr.$
This metric helps detect unphysical distortions like bond stretching or angle collapse, common causes of failure in MLFF-driven MD, providing a more interpretable and structure-specific complement to conventional force MAE.
@@ 2.3 Stability Criterion
We assess the stability of MD simulations for flexible molecules by monitoring bond lengths over time. A simulation is classified as unstable if, at any time $t<300\,\text{ps}$, the bond length between any atom pair exceeds a specified threshold, indicating bond breaking or excessive stretching that compromises structural integrity before the intended simulation duration. The criterion for instability is expressed as fu2023forces:
@@ 2.4 Machine Learning Force Fields
MLFFs learn an approximate representation of the potential energy surface from reference data, typically generated by quantum mechanical methods such as DFT. At each simulation step, the current atomic positions are input to the MLFF, which predicts the forces acting on each atom. These predicted forces are then used to iteratively update atomic positions throughout the simulation. Recently, GNN architectures have gained prominence for effectively capturing atomic connectivity in molecular systems. GNNs capture the connectivity of atomic structures by representing 3D atomic systems as graphs, where atoms are treated as nodes and interatomic distances define the edges gasteiger2021gemnet.
In this study, we utilize GemNet-T as the MLFF model. GemNet-T is a high-performing GNN designed to learn molecular energies and forces, building on prior models like DimeNet++ gasteiger2021gemnet. It enhances the representation of multi-scale geometric interactions by incorporating directional embeddings, capturing both radial and angular features. This enables more accurate force predictions, particularly for complex three-dimensional atomic arrangements.
Figure 1c illustrates the core structure of the GemNet-T model, a directional message-passing neural network designed for MD simulations. The model begins by embedding atomic information using both atomic numbers and local geometric features, including distances, angles, and dihedral angles. These inputs generate initial atom and edge embeddings that preserve the molecular structure. The embeddings are then processed through multiple stacked interaction blocks, each consisting of a directional message-passing module and an atom-wise update. In GemNet-T, the message passing operates over one-hop neighborhoods and uses triplet-based geometric information (from distance and angle inputs) to encode directional relationships. This approach avoids the complexity of higher-order interactions, such as the quadruplet-based dihedral terms used in GemNet-Q, while still capturing essential directional features gasteiger2021gemnet. After each message-passing step, atomic embeddings are updated through residual connections. Once all interaction layers are completed, the final atomic embeddings are aggregated to predict molecular energies and atomic forces. Importantly, these predictions are designed to be equivariant to translations, rotations, and reflections, ensuring that the model produces physically consistent results during MD simulations gasteiger2021gemnet.
@@ 2.5 Datasets
We use the MD17 dataset as the primary benchmark for small-molecule systems. MD17 consists of ab initio MD trajectories, where each snapshot provides atomic coordinates, DFT total energy, and per-atom forces chmiela2017machine. Aspirin, a medium-sized organic molecule with 21 atoms and multiple rotatable bonds, is selected as a representative system due to its structural flexibility. This flexibility amplifies the impact of small force prediction errors, potentially leading to unstable trajectories over time. The aspirin subset in MD17 includes over 20,000 configurations (i.e., molecular snapshots along an MD trajectory, each with atomic positions, energy, and forces). For this study, we use 9,500 configurations for training, 500 for validation, and 10,000 as a separate test set.
To extend the model’s generalization beyond isolated small molecules, we incorporate the Open Catalyst 2020 (OC20) dataset for pre-training. OC20 contains 1.3 million relaxations of molecular adsorptions on surfaces, supported by over 260 million corresponding DFT calculations chanussot2021open. We specifically utilize a 2-million-sample subset from the Structure-to-Energy-and-Force (S2EF) task, which provides forces and energies for each relaxed configuration. Unlike MD17, OC20 data include periodic boundary conditions, more accurately reflecting realistic catalytic surfaces. By training on this larger, chemically diverse dataset, we aim to equip the model with robust atomic interaction representations before fine-tuning on the smaller MD17 aspirin set.
@@ 3 Experiments
Figure 1: Overview of the MLFF framework for MD simulations. a. GemNet-T is trained on molecular (MD17) and catalyst (OC20) datasets, then used as a force predictor in MD simulations. b. Atomic positions are iteratively updated based on forces predicted by the trained GemNet-T model. c. Simplified illustration of the GemNet-T architecture, which incorporates multi-hop geometric message passing and directional embeddings to model atomic interactions.
@@ 4.1 Force Metric Limitations
We evaluate the stability of the MD simulations to determine whether high force prediction accuracy necessarily guarantees more stable simulation results. The MAE in force prediction serves as the primary error metric during MLFF training, as it is the objective function the model is optimized for. To evaluate dynamical stability, we use the instability onset time ($t_{\mathrm{inst}}$), which measures how long the simulation preserves structural integrity before any bond deviates beyond a defined threshold from its reference length. A higher instability onset time, approaching the 300 ps reference from ab initio MD, indicates greater stability and closer agreement with the ground truth trajectory. The instability onset times shown in Figure 2b represent the average values obtained from three independent simulations initialized with different random seeds.
As shown in Figure 2a, the force MAE curve decreases smoothly as training progresses. However, the difference between the pre-trained and non-pre-trained cases is minimal with a $\text{RMS Deviation}\approx 0.421\ \text{meV/\AA }$ at Epoch 300. This result makes it difficult to distinguish their performances based solely on force MAE.
In contrast, the MD stability trends diverge from the force MAE curve. Although the force MAE clearly decreases with further training, MD stability does not necessarily improve in tandem. In some cases, instability occurs earlier at later training epochs, reflected by a reduced instability onset time, despite better force prediction accuracy. This trend is particularly pronounced in the non-pre-trained case, suggesting that the model may overfit to the training dataset and struggle to generalize to diverse configurations encountered during extended MD simulations. This observation aligns with the findings of Fu et al. fu2023forces, indicating that force metrics alone may not reliably reflect downstream performance in MD simulations.
Figure 2: Impact of pre-training on force accuracy and simulation stability. a. Force MAE (meV/Å) over epochs for the GemNet-T model trained without pre-training (MD17 only) versus with pre-training (MD17 pre-training followed by OC20 finetuning). b. Instability onset time over epochs. Error bars represent results from three independent simulations.
4.2 Stability Enhancement